OpenStudy (anonymous):

quadratic formula of 2x^2+6x+4=0

OpenStudy (anonymous):

2 (x^2 + 3x + 2) = 0 x^2 + 3x + 2 = 0 (x + 2) ( x + 1) = 0 x = -2, -1